# Solving Optimal Power Flow on a Data-Budget: Feature Selection on Smart Meter Data

### v1

@@ VI Conclusions
We have proposed an OPF data distillation framework to improve the fidelity of either OPF features or OPF minimizers. Type-2 OPF data distillation gives rise to a non-convex program, solved by a proximal gradient algorithm. To alleviate the computational burden of finding the minimizers and their sensitivities for a batch of OPF instances, we have leveraged results from multiparametric programming. Extensive numerical tests using synthetic data on a single-phase and real-world data on a multiphase feeder demonstrate the advantages of Type-2 over Type-1 methods in approximating the minimizer and satisfying network constraints at the expense of increased complexity. Numerical tests have shown that optimal DER schedules can be approximated reasonably well upon collecting only $K=16$ rather than $P=50$ OPF features for the IEEE 37-bus single-phase feeder. For the multiphase system, DER schedules can be approximated within 25% error with $K=53$ out of $P=190$ features while nearly matching the voltage distribution of the exact OPF solutions. The proposed framework sets a solid foundation for exciting research directions: i) Extending data distillation to AC-OPF formulations or other OPF variants; ii Pursuing stochastic variants of APGD that process one or a few OPF instances per iteration to reduce the computational cost; iii) Leveraging OPF data distillation tools towards optimal meter placement to support OPF operations; iv) Upgrading the linear to a nonlinear reconstruction module (deep neural network); v) Pursuing unsupervised counterparts of OPF data distillation to waive the need of generating a labeled OPF dataset; and vi) Identifying the most influential OPF features so their communication to the DSO is protected against cyber-attacks.
@@
for all $p$. Function $g(\mathbf{W})$ is differentiable with respect to $\mathbf{w}_{p}$ for all $\mathbf{w}_{p}\neq\mathbf{0}$. Its subdifferential at $\mathbf{w}_{p}=\mathbf{0}$ can be defined as the set of all vectors having $\ell_{2}$-norm not larger than one.

### v2

@@ VI Conclusions
We have proposed an OPF data distillation framework to improve the fidelity of either OPF features or OPF minimizers. Type-2 OPF data distillation yields a non-convex program, which is solved by a proximal gradient algorithm. To alleviate the computational burden of finding the minimizers and their sensitivities for a batch of OPF instances, we have leveraged results from multiparametric programming.
Extensive numerical tests using real-world data on 37-bus and 1,136-bus single-phase feeders demonstrate the advantages of Type-2 over Type-1 methods in approximating the minimizer and satisfying network constraints at the expense of increased complexity. Numerical tests have shown that optimal DER schedules can be approximated reasonably well upon collecting only $K=16$ rather than $P=50$ OPF features for the IEEE 37-bus single-phase feeder. For the larger system, DER schedules can be approximated within 5% error using $K=116$ out of $P=1,038$ features while nearly matching the voltage distribution of the exact OPF solutions. For larger values of compression ratios $K/P$, the advantage of Type-2 methods diminishes in general, and Type-1 methods perform reasonably well.
The proposed framework sets a solid foundation for exciting research directions: i) Extending data distillation to AC-OPF formulations or other OPF variants; ii Pursuing stochastic variants of APGD that process one or a few OPF instances per iteration to reduce the computational cost; iii) Leveraging OPF data distillation tools towards optimal meter placement to support OPF operations; iv) Upgrading the linear to a nonlinear reconstruction module (deep neural network); v) Pursuing unsupervised counterparts of OPF data distillation to waive the need of generating a labeled OPF dataset; and vi) Identifying the most influential OPF features so that their communication to the DSO is protected against cyberattacks.
@@
for all $p$. Function $g(\mathbf{W})$ is differentiable with respect to $\mathbf{w}_{p}$ for all $\mathbf{w}_{p}\neq\mathbf{0}$. Its subdifferential at $\mathbf{w}_{p}=\mathbf{0}$ can be defined as the set of all vectors with an $\ell_{2}$-norm not exceeding one.